# OmniCT: Towards a Unified Slice-Volume LVLM for Comprehensive CT Analysis

### v1

@@ H.1 MedEval-CT-Bench
Across both MedEval-CT-Bench-2D (Table 9) and MedEval-CT-Bench-3D (Table 10), OmniCT consistently achieves the highest overall performance, with averages of 79.80 and 77.63, respectively, surpassing strong baselines such as GPT-5-mini (OpenAI, 2025), Lingshu, CT-CHAT, and M3D-LaMed. It demonstrates robust gains across diverse organs (e.g., liver, kidneys, heart, spine) and task levels (GIR, MAI, AII, CRD), excelling particularly in advanced interpretation and reasoning. These results highlight the effectiveness of our unified slice–volume paradigm in delivering stable, cross-task generalization and comprehensive CT understanding.
@@
(iii) Analysis of Organ-level Semantic Enhancement. The OSE module leverages organ segmentation as a structural regional prior rather than a supervision target. The segmentation masks indicate organ regions with high semantic load for typical CT-based reasoning, from which OSE aggregates a compact set of discriminative tokens, while all global tokens are preserved in the feature stream. In this way, OSE explicitly strengthens organ-level semantics without sacrificing global context. Since the module relies on organ-level structural consistency instead of pixel-level boundary fitting, the high stability of TotalSegmentor in thoracoabdominal CT (average Dice 94.3% (Wasserthal et al., 2023)) is well suited for providing such regional cues. To evaluate the effectiveness of OSE and to rule out potential bias introduced by the segmentation model, we designed three alternative strategies: (i) removing ROI regions, (ii) random ROI pooling, and (iii) directly concatenating native ROI tokens. As shown in Table 12, removing ROIs yields the expected performance drop; random pooling brings limited gains mainly due to weak alignment effects arising from repeated visual tokens; and direct concatenation of native ROI tokens produces variable-length sequences that prevent stable semantic compression and offer no performance benefit. In contrast, OSE’s fixed-dimensional adaptive aggregation preserves global information coverage while emphasizing diagnostically critical regions, making it better suited to the structured requirements of medical image analysis.
@@ H.3 Case Study
In Figure  8, the Ground Truth highlights portal hypertension, cirrhotic changes, intrahepatic bile duct cystic dilatations with the central dot sign, and additional renal and gallstones. M3D-LaMed-7B, while mentioning hepatic and portal abnormalities, overemphasizes tumor invasion with incorrect findings, and GPT-5 completely deviates with irrelevant head CT descriptions. In contrast, OmniCT accurately captures the key clinical features—portal hypertension, splenomegaly, mild ascites, cirrhosis, Caroli disease signs, and urinary/gallstones—closely matching the Ground Truth with clinically coherent language, demonstrating its superior spatial–semantic consistency in chest–abdominal CT interpretation.
@@
[…]
@@
[…]
@@
[…]
@@
[…]

### v2

@@ H.1 MedEval-CT-Bench
Across both MedEval-CT-Bench-2D (Table 9) and MedEval-CT-Bench-3D (Table 10), OmniCT consistently achieves the highest overall performance, with averages of 79.80 and 77.63, respectively, surpassing strong baselines such as GPT-5-mini (openai_gpt5_2025), Lingshu, CT-CHAT, and M3D-LaMed. It demonstrates robust gains across diverse organs (e.g., liver, kidneys, heart, spine) and task levels (GIR, MAI, AII, CRD), excelling particularly in advanced interpretation and reasoning. These results highlight the effectiveness of our unified slice–volume paradigm in delivering stable, cross-task generalization and comprehensive CT understanding.
@@
(iii) Analysis of Organ-level Semantic Enhancement. The OSE module leverages organ segmentation as a structural regional prior rather than a supervision target. The segmentation masks indicate organ regions with high semantic load for typical CT-based reasoning, from which OSE aggregates a compact set of discriminative tokens, while all global tokens are preserved in the feature stream. In this way, OSE explicitly strengthens organ-level semantics without sacrificing global context. Since the module relies on organ-level structural consistency instead of pixel-level boundary fitting, the high stability of TotalSegmentor in thoracoabdominal CT (average Dice 94.3% (wasserthal2023totalsegmentator)) is well suited for providing such regional cues. To evaluate the effectiveness of OSE and to rule out potential bias introduced by the segmentation model, we designed three alternative strategies: (i) removing ROI regions, (ii) random ROI pooling, and (iii) directly concatenating native ROI tokens. As shown in Table 12, removing ROIs yields the expected performance drop; random pooling brings limited gains mainly due to weak alignment effects arising from repeated visual tokens; and direct concatenation of native ROI tokens produces variable-length sequences that prevent stable semantic compression and offer no performance benefit. In contrast, OSE’s fixed-dimensional adaptive aggregation preserves global information coverage while emphasizing diagnostically critical regions, making it better suited to the structured requirements of medical image analysis.
@@ H.3 Case Study
In Figure 8, the Ground Truth highlights portal hypertension, cirrhotic changes, intrahepatic bile duct cystic dilatations with the central dot sign, and additional renal and gallstones. M3D-LaMed-7B, while mentioning hepatic and portal abnormalities, overemphasizes tumor invasion with incorrect findings, and GPT-5 completely deviates with irrelevant head CT descriptions. In contrast, OmniCT accurately captures the key clinical features—portal hypertension, splenomegaly, mild ascites, cirrhosis, Caroli disease signs, and urinary/gallstones—closely matching the Ground Truth with clinically coherent language, demonstrating its superior spatial–semantic consistency in chest–abdominal CT interpretation.
@@
Model Qwen2.5-VL InternVL3-8B RadFM Lingshu HealthGPT MedVLM-R1 MedGemma-4B GPT-5-mini Ours
@@
Model Qwen2.5-VL MiniCPM-V-4.5 CT-CHAT M3D-LaMed-Phi-3-4B M3D-LaMed-Llama-2-7B GPT-5-mini Ours
@@
Organ 2D Normal 2D Mismatch 2D Noise 3D Normal 3D Mismatch 3D Noise
@@
Training Strategy Ratio SLAKE VQA-RAD RadFig-VQA M3D-VQA CT-RATE 3D-RAD